OpenStudy (dominirican1013):

solve on the interval [0,2pi): 1+sin theta =1/2

OpenStudy (welshfella):

sin theta = 1/2 - 1 = -1/2 first use your calculator to find the angle whose sine is 1/2 -then you can find theta - recall that the sine is negative in the 3rd and 4th quadrants

OpenStudy (dominirican1013):

0.008727? or 30

OpenStudy (welshfella):

for sin-1 1/2

OpenStudy (welshfella):

30 degrees or pi/6

OpenStudy (welshfella):

i think they want the answer in radians as they wrote (0,2pi)

OpenStudy (dominirican1013):

yes

OpenStudy (welshfella):

but pi/6 is not the answer as the sine is negative 1/2 there are 2 solutions - one in 3rd and one in 4th quadrant

OpenStudy (welshfella):